dr/dΩ+tanΩ=cos2Ω


1
Expert's answer
2021-07-27T09:05:30-0400

Solution;

"\frac{dr}{d\Omega}" =cos2"\Omega" -tan"\Omega"

dr=(cos2"\Omega" -tan"\Omega" )d"\Omega"

Integrate both sides;

"\int{dr}=\int{(cos2\Omega-tan\Omega)}d\Omega"

"\int{dr}=\int{cos2\Omega}d\Omega-\int{tan\Omega}d\Omega"

"\int dr=\int{cos2\Omega}d\Omega-\int{\frac{sin\Omega}{cos\Omega}}d\Omega"

But for the integration of;

"\int{\frac{sin\Omega}{cos\Omega}}"

take u=cos"\Omega" ;"\frac{du}{d\Omega}=-sin\Omega"

"d\Omega=\frac{-1}{sin\Omega}du"

Replace back

"\int{\frac{-1}{u}}du=-ln(u)=-ln(cos\Omega)" =ln"(sec\Omega)"

Hence,

"r=\frac{sin2\Omega}{2}-ln(|sec\Omega|)+C"